OpenStudy (anonymous):

I cant get his problem, I'm working with Solving Quadratic Equations: the Principle of Square Roots 2x^(2)-8x=14 I know that i'm supposed to divide the entire question by 2, then take half of the coefficient of x, but its not coming out. What am i doing wrong?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

\[2x ^{2}-8x=14\]

OpenStudy (anonymous):

taking half throughout the equation u get, x^2-4x-7=0 then, x^2 -4x+4-4-7=0 => (x-2)^2=11 therefore we get, x=2+-sqrt(11);

OpenStudy (anonymous):

yan po ung sgot ?

OpenStudy (anonymous):

err...which language is that ?

OpenStudy (anonymous):

is that the answer ?

OpenStudy (anonymous):

yes that is the answer, let me look at what you did

OpenStudy (anonymous):

if u hav doubt post it :)

OpenStudy (anonymous):

do u all have a facebook ?

OpenStudy (anonymous):

\[2x ^{2}-8x=14\] then i moved the 14 over \[2x ^{2}-8x-14=0\] then i divided the entire equation by 2 \[x ^{2}-4x-7=0\] then i moved the 7 back over \[x ^{2}-4x=7\] so if i take the half of the coeffient of x, i end up with \[x2-4x+4=7\] i dont understand how you got \[\sqrt{11}\]

OpenStudy (anonymous):

this x^2 -4x+4-4-7=0 i dont understand

OpenStudy (anonymous):

use the quadratic formula

OpenStudy (anonymous):

ok let me try that

OpenStudy (anonymous):

in deriving the last equation , u forgot that wen u add 4 u need to subtract it too! get it?

OpenStudy (anonymous):

ok, i got it using the quadratic formula

OpenStudy (anonymous):

ur saying when i added for, i then need to subtract if from 7?

OpenStudy (anonymous):

four

OpenStudy (anonymous):

u hv to carry out same operation on both sides to maintain equality

OpenStudy (anonymous):

ok look at the last but one equation. u hav x^2-4x=7 and then u hav got x^2-4x+4=7 , which ur telling ,are the same 2 equations. wat i'm telling is, u should hav done x^2-4x+4=7+4

OpenStudy (anonymous):

ohh yes thats right

OpenStudy (anonymous):

but i still have x2-4x+4=11

OpenStudy (anonymous):

the x2-4x+4 cannot be factored

OpenStudy (anonymous):

oh pellet, i got it yes it can

OpenStudy (anonymous):

so it shud b x^2 - 4x +4 = 11 the left hand side become (x -2)^2 so (x-2)^2 = 11 so x+2 = +/- root11 x = -2 +/-root11

OpenStudy (anonymous):

\[(x-2)^{2}=\sqrt{11}\]

OpenStudy (anonymous):

yes, thank you all! for some reason i just could not see this

OpenStudy (anonymous):

no plus-minus root11

OpenStudy (anonymous):

yes the final is \[x=2+-\sqrt{11}\]

OpenStudy (anonymous):

\[(x-2) = \pm \sqrt{11}\]

OpenStudy (anonymous):

yes final is what u wrote

OpenStudy (anonymous):

thanks you! you all are great!

OpenStudy (anonymous):

i just could not get what i was missing

OpenStudy (anonymous):

its so silly when i look at it now

OpenStudy (anonymous):

Got a medal for US !!! click GOOD ANSWER button for all of us

OpenStudy (anonymous):

( :))

OpenStudy (anonymous):

if i find that button i will..lol

OpenStudy (anonymous):

just above a bar appears with my name Look to the extreme right of this bar n u will see the GOOD ANSWER button

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

tku very much !!!!